16. Two bodies, one of mass 4 slugs and the other of mass m is fastened to the two ends of a string. The string passes over a smooth pulley so that the two bodies hang vertically and the 4 slug mass is 4 ft from the floor. One second after the bodies are released, the 4 slug mass reaches the floor. Find the mass of the other body. 


1
Expert's answer
2021-09-27T09:03:25-0400

Find the acceleration from time and displacement:


"a=\frac{2d}{t^2}."

Find the acceleration from equations by NSL for both bodies:


"-m_1a=-m_1g+T,\\
m_2a=-m_2g+T.\\\space\\
a(m_1+m_2)=(m_1-m_2)g,\\\space\\

m_2=m_1\frac{g-a}{g+a}."

Substitute acceleration (the first eqn above):


"m_2=m_1\frac{g-2d/t^2}{g+2d/t^2}=2.4\text{ slug}."